# Arc-length characterization of finite, radial growth patterns

### v1

@@ Abstract
We present a method to characterize the distribution of length-scales of finite, disordered patterns with, on average, radial symmetry. This method makes it possible to quantify the distribution of characteristic length scales in cases where the conventional “linear” chord method does not work. We show that the method can clearly distinguish regular patterns, patterns that are formed by diffusion-limited aggregation, and patterns that form during the slow drying of confined, colloid-laden droplets, explained by Beechey-Newman et al. 2. We also introduce a method to find the centre-point of these finite patterns, without assuming a full connectivity in the pattern. The method should be widely applicable to other, finite quasi-two-dimensional patterns.
@@ 1 Introduction
The so-called “chord-length distribution method” 10 is a convenient tool to characterize the distribution of length scales of clusters and voids in a disordered yet homogeneous, two-dimensional pattern of two phases (e.g. solid and air). This chord analysis aims to determine the distributions of segment-lengths of a (any) straight line that intersects a large number of solids and void spaces, e.g. in a gel network 13. The conventional chord analysis does not provide meaningful results for patterns that are not homogeneous and finite. Yet such patterns are important: many quasi-two-dimensional aggregation phenomena yield patterns that are a) finite and b) not homogeneous in space. One example among many is the pattern that is formed by diffusion limited aggregation (DLA). For such, on average radial patterns, chord analysis is not useful. Here we propose a closely related approach, to probe the arc-length distribution of circles that intersect both solid and void domains. Importantly, this arc-length distribution depends on the radius of the intersecting circle, even when averaged over many similar patterns. In what follows, we will refer to this type of analysis as circular chord-length analysis: this name is less precise than arc-length distribution, but stresses the analogy with the conventional chord-length distribution. Arc-length distributions can serve as a “fingerprint” of disordered radial patterns, obtained experimentally as shown in Figure 1. We compare the circular arc-length distribution for three visually different patterns in Figure 3 (a colloidal drying pattern, a diffusion limited aggregation (DLA) pattern, and a “random birthday pie” pattern) to illustrate that the method can be used to quantify the visual differences between these patterns.
@@ 2 Systems
Roughly radial patterns often form due to the deposition of colloids, either due to evaporation of the solvent, or due to (diffusion-limited or reaction-limited) aggregation of colloids from a colloid-laden suspension. A classic example is the famous “coffee-stain-effect” 3 where the dispersed coffee grounds mostly accumulate in the rim of the dried stain. The effect was explained in ref. 3 for the case of droplets on partially wetting surfaces, where the droplet circumference gets pinned. As the droplet evaporates, there will be a capillary flow towards the rim, transporting almost all colloidal particles towards the perimeter. In inkjet printing, but also in coatings and paintings, this effect is unwanted, and efforts have been made to overcome the capillary flows that drive deposition on the pinned contact line 6, 1, 9, 8.
@@
The fractal dimension of a pattern is a convenient quantity to characterize naturally-formed radial patterns that are scale-invariant. This fractal dimension can be used to distinguish different growth mechanisms; e.g., viscous fingering, stable displacement or capillary fingers in the invasion percolation of two immiscible fluids in porous media 4. In contrast, the patterns observed in ref. 2 are not  scale-invariant: the characteristic length scale of the branches changes rather abruptly with radius, something that does not lend itself for description in terms of a fractal or multi-fractal description. It is for this reason, that for such systems, we explore the angular equivalent of the chord method.
The remainder of this paper is organized as follows: first we introduce the circular arc-length distribution method (CAL), and its relation to the linear chord-length distribution introduced by Méring and Tchoubar 10. We then consider examples of applications of the CAL method to experimentally observed drying patterns of colloidal droplets, for DLA patterns, and for and artificial radial patterns. We show that, for experimental patterns, the CAL method can generate unique fingerprints that depend on the radius of the intersecting circle. Such an analysis is particularly useful in cases where the nature of the pattern formation changes with the radius of the deposit.
@@ 3 Arc-length distributions
The circular arc-length analysis used in this work is a fairly straightforward generalization of the linear chord-length method of Levitz and Tchoubar 7. In its simplest form, the CAL method is used to characterize biphasic patterns, where the two “phases” might be real phases of matter, e.g., solid and vapor, or simply different labels (e.g., black and white, characterizing an artificially generated pattern).
We use a fairly standard method to binarize the experimental images, similar to the method used in Zupkauskas et al. 13. First, we apply a Gaussian blur on the image, and threshold by Otsu’s method 11. Finally, a labelling algorithm is used to filter out all clusters below a certain size, to minimize the random labelling errors that inevitably occur.
@@
Once the centre of the pattern has been located or chosen, we can draw circles of increasing radius $r$ around this centre. Typically, such a circle will intersect many boundaries between the two phases (say: “solid” and “void’). If we attribute the value $1$ to all points in the solid phase, and 0 to points in the void phase, then we can construct a function $\mathcal{T}(\theta,r)$, where $0<\theta\leq 2\pi$ is the angle specifying a point on the circle corresponding to radius $r$. As $\mathcal{T}(\theta,r)$ can have only the two values 0 and $1$, the resulting function at constant $r$ looks like a (random) telegraph signal as a function of $\theta$. In experiments, the resolution with which the pattern is obtained is limited by the pixel size, typically a square with length and width $\Delta$. A pixel “belongs” to a given circle if that circle intersects the pixel area. In general, a pixel might belong to more than one circle. However, if the radii of adjacent circles differ by more than $\Delta\sqrt{2}$ then every pixel belongs to at most one circle. Of course, the density of pixels on a circle will, in general, not be constant, but in practice that is not a serious problem for determining arc-segment lengths.
Figure 2: (a) Illustration of the circular arc-length distribution for the colloidal drying pattern P2, that was appropriately cropped and binarized. For a given radius (red circle), each void-solid (white-black) interface is marked with a red dot, allowing us to store the length along the circle between each interface. (b) Resulting distribution of the arc lengths $s$ between interfaces, where $p(s)$ is the probability for a given $s$. Radii lying in the orange shell show a broader distribution than those lying in the blue, inner shell, with finer structures. (c) Telegraph signal $\mathcal{T}(\theta)$ obtained from the particular radius illustrated in (a), with $\theta$ starting from the 3 o’ clock position moving clockwise. In the signal, 1 indicate the solid phase and 0 corresponds to the void phase. The details of the CAL distribution differ from sample-to-sample, but the overall characteristics remain the same.
@@ 4 Analysis of experimental and synthetic patterns
We tested the CAL method by applying it to the the quasi-2D colloidal drying patterns P1, P2 and P3 in Figure 3. These patterns formed under the conditions explained in Figure 1, which differ qualitatively. In addition, we applied the same circular-arc analysis to a numerically generated two-dimensional diffusion-limited aggregation model and an artificial pattern of random segments, which we refer to as the “random birthday pie” (RBP). As was discussed by Levitz and Tchoubar 7, DLA patterns yield a power-law distribution of the chord lengths in the void phase: $p_{p}(s)\sim 1/s^{d_{m}-1}$ where $s$ is the chord length and $d_{m}$ is the fractal dimension of the pattern in the void phase. Below, we first apply the circular arc-length analysis to a radial colloidal drying pattern 2 and then compare it to that of a DLA pattern. As the mechanisms by which the drying and DLA patterns form are very different, we might expect that this should result in rather different CAL distributions.
@@
In the conventional chord length analysis, where we create histograms of cord-lengths in $x$- and $y$-directions in a porous medium, for instance, we are usually interested in finding an average characteristic length scale. This could be for instance the average pore-size in the system 13. In our radial analogy, we illustrate the mean arc lengths For the experimentally obtained and artificial patterns in Figure 4. The RBP pattern scales linearly in both phases (with the solid fraction $\varphi\approx 0.5$ as the slope; turquoise dashed lines). In contrast, the DLA pattern has a constant solid arc length, while the average void arc-length grows as a power-law with the radius (black dashed lines). We find that the drying patterns fall somewhere between the two synthetic patterns. All colloidal drying patterns have on average thicker fingers (solid phase) and thicker voids (void phase) at the outer radii. The main difference is that the finger thickness and void thickness are comparable close to the centre, but the average void finger thickness grow much faster.
Figure 4: Average arc-lengths of both phases, averaged over $dr=20$ pixels, for the 3 drying patterns P1-P3. The growth trends of the synthetic patterns are added as dashed lines for illustration.
We can also explore other properties of the telegraph signal we find in our radial chord length analysis. Starting with the smallest binning length of $dr=1$ pixel, we study two properties with respect to the radius. In Figure 5, we present the fraction of the solid phase (i.e. by counting the frequency of 1 in the telegraph signal) of the drying patterns at each radius. We expect $\varphi=0$ far away where we have a monolayer deposition and therefore do not have any clear fingers 2; likewise we expect that the centre of the pattern should be in the solid phase, and expect $\varphi=1$ as $r\rightarrow 0$. All 3 colloidal maze patterns studied display 3 different characteristics in $\varphi(r)$: P1 oscillates around a stable value, where the peaks in the oscillations correspond to the radii at which branching occurs; in pattern P2 the solid fraction increases as the pattern dries; pattern P3 has two fairly sharp jumps with a stable value between them. As these patterns are formed in similar conditions, the instability offer a surprising sensitivity to the subtle differences in the initial conditions.
When inspecting the fingers in a microscope, it is clear that they are fairly uniform in colloid density across the pattern, whereas the space between the fingers contain no colloids. Therefore, measuring the solid and void phases in these images will, to a fairly good approximation, be the same as measuring the mass of the colloidal deposition.
@@
At each radius, it is possible to count the number of discontinuous jumps in the telegraph signal. Each of these jumps represents a change in the phase (i.e. an interface); the length between phase changes is the arc lengths we study. Thus, counting the number of jumps in the signal is the same as counting the number of distinct arc lengths, and is therefore proportional to the number of fingers, at each radius. The scaling is known for the DLA and Birthday Pie: DLA creates fingers linearly with its growth, while a radially invariant pattern has the same number of interfaces at every radius. The phase jumps of our drying patterns are shown in Figure 6 and display a transition between the characteristic behavior of DLA, close to the centre, and RBP far away from the centre. The fingers clearly merge as the droplet dries, while the fluid instability causing the fingers also form new fingers during drying. The scaling in region (III) in Figure 6 means the rate of merging roughly matches the rate of forming new fingers. Region (II) marks the transition, where the fingers merge more frequently, and finally, in region (I) the colloidal fingers follow a linear ”DLA-like” decay towards the centre.
@@
Figure 6: Number of phase shifts of the telegraph signal, corresponding to the surface area the fingers generate. In a DLA, the number of interfaces grows with the circumference, while radially invariant patterns, e.g. Random Birthday Pies, have a constant number of phase shifts. We mark this figure with 3 regions where the colloidal patterns exhibit different scaling: Region (III) is characterised with a fairly stable number of interfaces, in region (II) the fingers start merging faster than they are formed, and in region (I) there is a fairly stable merging rate across the drying patterns.

### v2

@@ Abstract
We present a method to characterize the distribution of length-scales of finite, disordered patterns with, on average, radial symmetry. This method makes it possible to quantify the distribution of characteristic length scales in cases where the conventional “linear” chord method does not work. We show that the method can clearly distinguish regular patterns, patterns that are formed by diffusion-limited aggregation, and patterns that form during the slow drying of confined, colloid-laden droplets, explained by Beechey-Newman et al. 2. We also introduce a method to find the centre-point of these finite patterns, without assuming a full connectivity in the pattern. The method should be widely applicable to other, finite quasi-two-dimensional patterns like dendritic structures, viscous fingering, liquid crystal patterns and bacterial growth.
@@ 1 Introduction
The so-called “chord-length distribution method” 13 is a convenient tool to characterize the distribution of length scales of clusters and voids in a disordered yet homogeneous, two-dimensional pattern of two phases (e.g. solid and air). This chord analysis aims to determine the distributions of segment-lengths of a (any) straight line that intersects a large number of solids and void spaces, e.g. in a gel network 18. The conventional chord analysis does not provide meaningful results for patterns that are not homogeneous and finite. Yet such patterns are important: many quasi-two-dimensional aggregation phenomena yield patterns that are i) finite and ii) not homogeneous in space. One example among many is the pattern that is formed by diffusion limited aggregation (DLA). For such, on average radial patterns, chord analysis does not provide a meaningful description. Here we propose a closely related approach, to probe the arc-length distribution of circles that intersect both solid and void domains. Importantly, this arc-length distribution depends on the radius of the intersecting circle, even when averaged over many similar patterns. In what follows, we will refer to this type of analysis as circular chord-length analysis: this name is less precise than arc-length distribution, but stresses the analogy with the conventional chord-length distribution. Arc-length distributions can serve as a “fingerprint” of disordered radial patterns, obtained experimentally as shown in Figure 1. We compare the circular arc-length distribution for three visually different patterns (3 typical colloidal drying patterns; a diffusion limited aggregation (DLA) pattern; and a “random birthday pie” pattern) to illustrate that the method can be used to quantify the visual differences between these patterns.
@@ 2 Systems
Roughly radial patterns often form due to the deposition of colloids, either due to evaporation of the solvent, or due to (diffusion-limited or reaction-limited) aggregation of colloids from a colloid-laden suspension. A classic example is the famous “coffee-stain-effect” 3 where the dispersed coffee grounds mostly accumulate in the rim of the dried stain. The effect was explained in ref. 3 for the case of droplets on partially wetting surfaces, where the droplet circumference gets pinned. As the droplet evaporates, there will be a capillary flow of the solvent towards the rim, transporting almost all colloidal particles towards the perimeter. In inkjet printing, but also in coatings and paintings, this effect is unwanted, and efforts have been made to overcome the capillary flows that drive deposition on the pinned contact line 8, 1, 12, 11.
@@
The fractal dimension of a pattern is a convenient quantity to characterize naturally-formed radial patterns that are scale-invariant. This fractal dimension can be used to distinguish different growth mechanisms; e.g., viscous fingering, stable displacement or capillary fingers in the invasion percolation of two immiscible fluids in porous media 4. In contrast, the patterns observed in ref. 2 are not  scale-invariant: the characteristic length scale of the branches changes rather abruptly with radius, something that does not lend itself to a fractal or multi-fractal description. It is for this reason, that for such systems, we explore the angular equivalent of the chord method.
The remainder of this paper is organized as follows: first we introduce the circular arc-length distribution method (CAL), and its relation to the linear chord-length distribution introduced by Méring and Tchoubar 13. We then consider examples of applications of the CAL method to experimentally observed drying patterns of colloidal droplets, for DLA patterns, and for and artificial radial patterns. We show that, for experimental patterns, the CAL method can generate unique fingerprints that depend on the radius of the intersecting circle. Such an analysis is particularly useful in cases where the nature of the pattern formation changes with the radius of the deposit.
@@ 3 Arc-length distributions
The circular arc-length analysis used in this work is a fairly straightforward generalization of the linear chord-length method of Levitz and Tchoubar 10. In its simplest form, the CAL method is used to characterize biphasic patterns, where the two “phases” might be real phases of matter, e.g., solid and vapor, or simply different labels (e.g., black and white, characterizing an artificially generated pattern), as seen in Figure 2.
We use a fairly standard method to binarize the experimental images, similar to the method used in Zupkauskas et al. 18. First, we apply a Gaussian blur on the image, and the threshold it using Otsu’s method 14. Finally, a labelling algorithm is used to filter out all clusters below a certain size, to minimize the random labelling errors that inevitably occur.
@@
Once the centre of the pattern has been located or chosen, we can draw circles of increasing radius $r$ around this centre. Typically, such a circle will intersect many boundaries between the two phases (say: “solid” and “void’). If we attribute the value $1$ to all points in the solid phase, and 0 to points in the void phase, then we can construct a function $\mathcal{T}(\theta,r)$, where $0<\theta\leq 2\pi$ is the angle specifying a point on the circle corresponding to radius $r$. As $\mathcal{T}(\theta,r)$ can have only the two values 0 and $1$, the resulting function at constant $r$ looks like a (random) telegraph signal that depends on $\theta$. In experiments, the resolution with which the pattern is obtained is limited by the pixel size, typically a square with length and width $\Delta$. A pixel “belongs” to a given circle if that circle intersects the pixel area. In general, a pixel might belong to more than one circle. However, if the radii of adjacent circles differ by more than $\Delta\sqrt{2}$ then every pixel belongs to at most one circle. Of course, the density of pixels on a circle will, in general, not be constant, but in practice that is not a serious problem for determining arc-segment lengths.
Figure 2: (a) Illustration of the circular arc-length distribution for the colloidal drying pattern P2, that was appropriately cropped and binarized. For a given radius (red circle), each void-solid (white-black) interface is marked with a red dot, allowing us to store the length along the circle between each interface. (b) Resulting distribution of the arc lengths $s$ between interfaces, where $p(s)$ is the probability for a given $s$. Radii lying in the orange shell show a broader distribution than those lying in the blue, inner shell, with finer structures. (c) Telegraph signal $\mathcal{T}(\theta)$ obtained from the particular radius illustrated in (a), with $\theta$ starting from the 3 o’ clock position moving clockwise. In the signal, 1 indicates the solid phase and 0 corresponds to the void phase. The details of the CAL distribution differ from sample-to-sample, but the overall characteristics remain the same.
@@ 4 Analysis of experimental and synthetic patterns
We tested the CAL method by applying it to the quasi-2D colloidal drying patterns P1, P2 and P3 in Figure 3. Although these patterns were formed under similar conditions, they can differ qualitatively as explained by Beechy-Newman et al.2. In addition, we applied the same CAL analysis to a numerically generated, two-dimensional diffusion-limited aggregation model 17 and an artificial pattern of random segments, which we refer to as the “random birthday pie” (RBP). As was discussed by Levitz and Tchoubar 10, DLA patterns yield a power-law distribution of the arc lengths in the void phase: $p_{p}(s)\sim 1/s^{d_{m}-1}$ where $s$ is the arc length and $d_{m}$ is the fractal dimension of the pattern in the void phase. Below, we first apply the circular arc-length analysis to a radial colloidal drying pattern 2 and then compare it to that of a DLA pattern. As the mechanisms by which the drying and DLA patterns form are very different, we might expect that this should result in rather different CAL distributions.
@@
In the conventional chord length analysis, where we create histograms of cord-lengths in $x$- and $y$-directions in a porous medium, for instance, we are usually interested in finding an average characteristic length scale. This could be for instance the average pore-size in the system 18. In our radial analogy, we illustrate the mean arc lengths for the experimentally obtained and artificial patterns in Figure 4. The RBP pattern scales linearly in both phases (with the solid fraction $\varphi\approx 0.5$ as the slope; turquoise dashed lines). In contrast, the DLA pattern has a constant solid arc length, while the average void arc-length grows as a power-law with the radius (black dashed lines). We find that the drying patterns fall somewhere between the two synthetic patterns. All colloidal drying patterns have on average thicker fingers (solid phase) and thicker voids (void phase) at the outer radii. The main difference is that the finger thickness and void thickness are comparable close to the centre, but the average void finger thickness grow much faster.
Figure 4: Average arc-lengths of both phases, averaged over $\mathrm{d}r=20$ pixels, for the 3 drying patterns P1-P3. The growth trends of the synthetic patterns are added as dashed lines for illustration.
We can also explore other properties of the telegraph signal we find in our radial chord length analysis. Starting with the smallest binning length of $\mathrm{d}r=1$ pixel, we study three properties with respect to the radius: Area fractions, number of interfaces and cross-radius correlation functions. In Figure 5, we present the fraction of the solid phase (i.e. by counting the frequency of 1 in the telegraph signal) of the drying patterns at each radius. We expect $\varphi=0$ far away where we have a monolayer deposition and therefore do not have any clear fingers 2; likewise we expect that the centre of the pattern should be in the solid phase, and expect $\varphi=1$ as $r\rightarrow 0$. All 3 colloidal maze patterns studied display 3 different characteristics in $\varphi(r)$: P1 oscillates around a stable value, where the peaks in the oscillations correspond to the radii at which branching occurs; in pattern P2 the solid fraction increases as the pattern dries; pattern P3 has two fairly sharp jumps with a stable value between them. As these patterns are formed in similar conditions, the fluid instability offer a surprising sensitivity to the subtle differences in the initial conditions.
@@
At each radius, it is possible to count the number of discontinuous jumps in the telegraph signal. Each of these jumps represents a change in the phase (i.e. an interface); the lengths between phase changes are the arc lengths we study. Thus, counting the number of jumps in the signal is the same as counting the number of distinct arc lengths, and is therefore proportional to the number of fingers, at each radius. The scaling is known for the DLA and Birthday Pie: DLA creates interfaces as a power-law with its fractal dimension $N\sim r^{(D_{f}-1)}$9 while a radially invariant pattern has the same number of interfaces at every radius. The phase jumps of our drying patterns are shown in Figure 6 and display a transition between the characteristic behavior of DLA, close to the centre, and RBP far away from the centre. The fingers clearly merge as the droplet dries, while the fluid instability causing the fingers also form new fingers during drying. The scaling in region (II) in Figure 6 means the rate of merging roughly matches the rate of forming new fingers. At some point there is a transition where the fingers merge more frequently, leading to region (I), where the colloidal fingers follow a linear ”DLA-like” decay towards the centre.
@@
Figure 6: Number of phase shifts of the telegraph signal, corresponding to the surface area the fingers generate. In a DLA, the number of interfaces grows with the circumference, while radially invariant patterns, e.g. random birthday pies, have a constant number of phase shifts. We mark this figure with two regions where the colloidal patterns exhibit different scaling: Region (II) is characterised with a fairly stable number of interfaces; at some point, the fingers start merging faster than they are formed, and in region (I) there is a fairly stable merging rate across the drying patterns.
We may also study the correlation between telegraph signals, to quantify the structural similarity between radii. Here, we want to factor out the correlation between solid fractions $\varphi$; we therefore remove the average from the telegraph signal and only look at the fluctuating part $\widetilde{\mathcal{T}}(r,\theta)=\mathcal{T}(r,\theta)-\langle\mathcal{T}(r,\theta)\rangle_{\theta}$, with $\langle A(\theta\rangle_{\theta}=\frac{1}{2\pi}\int_{0}^{2\pi}A(\theta)\mathrm{d}\theta$. The cross-correlation function of the pattern is defined as
$C(r,r^{\prime})=\frac{\left\langle\widetilde{\mathcal{T}}(r,\theta)\widetilde{\mathcal{T}}(r^{\prime},\theta)\right\rangle_{\theta}}{\sqrt{\left\langle\widetilde{\mathcal{T}}(r,\theta)^{2}\right\rangle_{\theta}\left\langle\widetilde{\mathcal{T}}(r^{\prime},\theta)^{2}\right\rangle_{\theta}}}.$ (1)
A pattern that is radially invariant, such as the RBP patterns, will have the property that $\mathcal{T}(r,\theta)=\mathcal{T}(r^{\prime},\theta)$; these are perfectly correlated at all radii, with only a small numerical error due to rasterization of the image. The correlation for the DLA and P1-P3 patterns are shown in Figure 7. We can see that the drying patterns P1-P3 decorrelate faster close to the centre, and are more persistent in the outer regions, where the fingers initially form.
Figure 7: Telegraph signal correlation matrix for patterns P1-P3 and DLA.
We then use equation (1) to find the average correlation of the separation of the compared signals as $C(\Delta r)=\big\langle C(r,r^{\prime})\big\rangle_{|r-r^{\prime}|=\Delta r}$. We plot these in Figure 8, where the decorrelation of the 3 drying patterns collapse to the same exponential decay curve. We introduce the persistence length $\xi$ as the length where the signal has decorrelated by a factor $1/e$, i.e. $C(\Delta r)=\exp(-r/\xi)$.
Figure 8: Average correlations of telegraph signals, by separation distances $\Delta r$ for all the 5 patterns. The RBP remains perfectly correlated throughout the pattern, while the DLA decorrelates rapidly – with a correlation length of 2 pixels. The drying patterns P1-P3 decay exponentially, with decay lengths of 0.063, 0.065 and 0.061 mm. We plot the average decay length of $\xi=0.063$ $m$.
The persistence length is found to be $0.063\pm 0.002$ mm for the drying patterns, which translates to about 35 colloid diameters. Physically, this is how long the fingers tend to stay aligned, and is therefore a characteristic length between the branching points. Likewise, we may also find the correlation between angles in the telegraph signal, where the arc-length is the natural variable to compare the decorrelation. We find that these decorrelation lengths give similar results as the mean arc-lengths, shown in Figure 4.
5 Estimating the colloidal concentration from arc lengths
[…]
$m_{c}(r)=\rho_{TPM}\cdot 2\phi\cdot d_{c}\int_{0}^{r}\varphi(r^{\prime})\cdot 2\pi r^{\prime}\mathrm{d}r^{\prime}.$ (2)
This relation between the mass and radius of a pattern is useful for probing the fractality of the patterns. A fractal will have a power-law relation between the mass deposition, scaling with its fractal dimension $m_{d}(r)\sim r^{D_{f}}$ 4. In comparison, the scaling of the RBP is simply its Euclidean dimension $m_{d,RBP}=\varphi\pi r^{2}$. We find that the mass scaling of patterns P1-P3 are well described by power-law fits, meaning the patterns may be described as fractals, with a fractal dimension similar to that of the DLA pattern (see Figure 9). We also measure the fractal dimension with a box-counting method, all summarized in Table 1. These values are fairly similar to the fractal dimension of a DLA pattern 17, except for the mass fractal dimension of P1, which is closer to the one of a RBP.
Table 1: Comparison of the fractal dimensions of the patterns P1-P3, measured with two different methods. The mass scaling method is based on a power-law-fit to the mass deposition data, while the box-counting method measures how many boxes NN of size $\delta$ we need to cover the full pattern, with the fractal dimension found as $D_{f}=\ln{N}/\ln{\delta}$.
Mass scaling dimension Box-counting dimension
P1 1.95 1.61
P2 1.67 1.68
P3 1.69 1.72
[…]
Figure 9: A plot of the cumulative colloid mass $m_{c}(r)$ deposited in the pattern forming stage of the evaporation process. This is found by assuming a uniform colloidal packing fraction throughout the pattern, left in exactly 2 monolayers. We find that the mass deposition scales with a power-law, with exponents found in Table 1. We then relate the deposited mass to the colloid concentration of the droplet as it is evaporating $C(r)$, compared to the initial concentration $C_{0}=1$ wt% for the patterns P1-P3. The initial concentration of the droplet is illustrated with a dashed black line.
[…]